How much energy is lost by 20 kg copper if it is cooled from 50°C to 30°C?


1
Expert's answer
2021-05-03T10:28:00-0400

The specific heat capacity of copper is 0.385 J/goC, which equals 385 J/kgoC

"Q=cm\Delta{T}=385\frac{J}{kg\cdot{\degree{C}}}\times20kg\times(50\degree{C}-30\degree{C})=154000J=154kJ"


Answer: 154 kJ